Help please
Frequency Scores 50 to 59 6010 69 70 to 79 80 to 89 90 to 99 What is the class width? O A. 10 0B.9 ° C. correct answer is not given O D. it can not be determined If a cumulative frequency polygon was made of the data, what would be the cumulative frequency of the thrid class? OA, 10 O B. correct answer is not given O c. 12
Answer #1

Question 1)

Class width is 10

Class width is distance between (Lower /upper ) limits of consecutive classes.

For example: 60-50 = 10 , 70-60 = 10

Option A) is correct.

Question 2)

Cumulative frequency of third class is 2+4+6 = 12

Option C) is correct.